### v2

@@ 1 Introduction
The rest of the paper is organized as follows: First, we provide a brief review of relevant work in Section 2. In Section 3, we detail Cornell University’s final exam scheduling problem. Section 4 outlines our proposed solution methods, with detailed optimization models and algorithms provided in the Appendix. In Section 5, we present the outcomes of our scheduling models and algorithms using real student enrollment data from Cornell University. Finally, Section 6 concludes our work.
@@ 2.1 Educational Timetabling
Educational timetabling includes three main subfields: high-school timetabling, course timetabling, examination timetabling. Many real-world problems in these areas are approached through integer programming techniques. High-school timetabling, commonly represented in the XHSTT format (Post et al. 2016), has seen the application of MIP in studies like Kristiansen et al. (2015) and Fonseca et al. (2017). Course timetabling problems, such as those at the United States Air Force Academy (Gonzalez et al. 2018), the Technion in Israel (Strichman 2017), and MIT (Barnhart et al. 2022), also employ these approaches. In examination timetabling, there are case studies from the Universidad Politécnica de Madrid (García-Sánchez et al. 2019) and the Department of Education, Vestfold County (Avella et al. 2022)
@@ 2.3 University Examination Timetabling: Solution Techniques
2.4 Contributions
Our main contribution are as follows:
  1. 1.
  2. 2.
    We develop an integrated MIP-based framework that synergizes local search and heuristic techniques to efficiently generate exam schedules optimized according to multiple performance metrics. Furthermore, our framework supports hybridization strategies to further enhance solution quality.
  3. 3.
3 Problem Description
Cornell University faces the complex task of scheduling final exams for a significant student body of more than 15,000 and approximately 700 courses at the beginning of every semester. These exams are concentrated within a 7 to 9-day final exam period, with three time slots available per day (9am, 2pm, and 7pm). However, certain time slots—such as Saturday nights and Sunday mornings—may have restrictions, resulting in roughly 19 to 22 available time slots. The ultimate scheduling problem is to assign around 500 exams to these time slots. The physical locations for the exams are determined separately by the University Registrar, and as such, this study focuses solely on optimizing the timing of the exams.
The goal of an optimized final exam schedule is to enhance the experience for both students and faculty. From the faculty’s perspective, student conflicts—such as such as overlapping or consecutive exams–may require the scheduling of make-up exams according to university policies (https://registrar.cornell.edu/calendars-exams/final-exam-policies), which imposes additional burdens. Meanwhile, students desire a well-distributed exam schedule to allow for sufficient review and study time. As a result, the optimal exam schedule should minimize various unfavorable metrics, each with its own weight. These metrics include direct conflicts, back-to-back exams, two exams in 24 hours, three exams in a row, and three exams in 4 slots.
@@ 4.1 Multi-Stage Group-then-Sequence (GtS) Model
The idea behind the first two stages is to initially group the exams into a smaller number of exam blocks and then assign each block to a time slot. The full IP formulations and pseudocode of the algorithms are provided in Appendix 7 - 11.
@@ 4.1.1 Block Assignment
The Block Assignment stage is formulated as a MIP problem (Min_Conflict_IP) that assigns each exam to one of a specified number of blocks, denoted by $k$. It is assumed that $k\leq|S|$, where $S:=\{0,1,\ldots,|S|-1\}$ is the set of available time slots. The objective of this grouping is to minimize the total number of pairwise co-enrollments (direct conflicts) among exams within the same block. It is important to note that, at this stage, each block represents a group of exams and is not yet associated with a specific time slot.
@@
We also propose a variation of the Block Assignment problem, named Zero_Conflict_IP, that imposes a strict zero-conflict constraint. This model incorporates an objective inspired by the Traveling Salesman Problem (TSP), aiming to minimize pairwise co-enrollments between exams in neighboring blocks. The TSP objective is designed to incorporate the goal of reducing the number of back-to-back exams during the Block Assignment stage. Figure 2 presents an example of a zero-conflict block assignment alongside its corresponding TSP objective. Compared to the minimum-conflict version, the zero-conflict problem has a more constrained feasible space, which increases the likelihood of infeasible solutions.
The feasibility of having a zero-conflict block assignment depends mainly on the required number of blocks $k$. While one way to find a lower bound on $k$ is to enumerate all possible values and check the feasibility of the Zero_Conflict_IP problem, a smarter way is to solve for the Maximum Clique (Max_Clique_IP) problem. The Max_Clique_IP involves finding a subset of vertices in $G$ that forms a complete sub-graph, where every two distinct vertices are adjacent. In our context, the size of the max clique corresponds to a lower bound on the number of blocks for which it is feasible to have no conflicts.
Figure 2: This is an example of a zero-conflict block assignment. The value on each edge represents the number of co-enrollments. Blocks 1 and 2, as well as Blocks 2 and 3, are considered neighboring blocks. Consequently, the TSP objective value can be calculated as the sum of the highlighted values.
@@ 4.1.2 Block Sequencing
After assigning exams to the designated blocks, we then allocate these blocks to time slots by sequencing them in Block Sequencing stage (Block_Sequencing_IP). In the Block Assignment phase, we focus mainly on direct conflicts, but in Block Sequencing, we begin to account for other levels of undesirable metrics.
@@ 4.1.3 Post-Processing Local Search Algorithm
Multiple optimal solutions with the same number of direct conflicts can be generated in the Block Assignment stage. However, the schedules derived from these assignments may vary in other metrics, as the Block Assignment model only focuses on direct conflicts. As a result, the block assignment obtained may not be optimal when considering all metrics comprehensively. To address this limitation, we develop a local search algorithm that takes an existing exam schedule and iteratively reschedules subsets of exams with a specified size $n$.
In each iteration, the algorithm selects the top $n$ exams to reschedule by assessing the contribution of each exam to the number of bad metrics, normalized by the number of students taking the exam. For example, an exam where 30% of its students have an exam right before or after it would be prioritized for rescheduling compared to an exam with only 5% of students in a similar situation. Once the subset is selected, the algorithm solves a Schedule_IP problem to move the exams to time slots that minimize the number of B2Bs and 2-in-24hrs, without increasing the number of direct conflicts.
The process continues iteratively, with the algorithm using adaptive step sizes to adjust the rolling window of size $n$. If no improvement is found after rescheduling, a larger step forward is taken to explore new solutions. Conversely, if an improvement is observed, a smaller step backward is adopted to fine-tune the schedule. This adaptive approach enables the algorithm to efficiently navigate the solution space, balancing exploration with refinement.
@@ 4.2 Layer-Cake Heuristic Algorithm
The Layer-Cake heuristic algorithm offers an alternative to the GtS model and is inspired by the Post-Processing algorithm. Unlike the GtS model, Layer-Cake addresses conflicts and other metrics simultaneously by solving the Schedule_IP problem, which optimizes all metrics together. To reduce the computational complexity of scheduling all exams at once, the problem is solved for a smaller subset of exams each time. A key difference between Layer-Cake and GtS is that Layer-Cake directly takes in a sequence of time slots, eliminating the need to specify the number of blocks, $k$, as an input.
Figure 5: Visualization of the Layer-Cake Algorithm
Layer-Cake follows a layered approach, where each layer consists of a set of exams such that the total number of students per layer is similar. The layers are determined by sorting the exams in descending order of student enrollment. The initial layers focus primarily on larger exams, while subsequent layers include many smaller exams. The algorithm starts by scheduling the largest exams independently of the others. It then schedules the next subset of exams on top of those already assigned, gradually building the schedule “layer by layer,” as shown in Figure 5. This ordering ensures that the largest exams are scheduled first, as the Schedule_IP problem becomes more challenging when dealing with larger exams compared to smaller ones, even with a constant total number of students.
An essential feature of the Layer-Cake algorithm is the overlap between consecutive layers. In each iteration, exams that have already been scheduled are included in the IP, allowing them to be rescheduled alongside the newly introduced exams. This overlapping of layers, serving as a warm start for the IP problems, considerably boosts the algorithm’s performance.
@@ 5 Implementation Results and Impact
Once the team gains access to the necessary data, we conduct experiments using various parameters and models implemented in Gurobi. This collaborative approach allows for comprehensive exploration and evaluation of different scheduling options. To accelerate the process, we use Google Colab to run multiple experiments in parallel. From the generated subset of “good” schedules, the University Registrar selects the best one to be published on the university website.
In the subsequent parts, we showcase the performance of our scheduling models on real student enrollment data, highlighting the framework’s ability to outperform the historical baseline and its flexibility in meeting the registrar’s needs.
It is important to note that while the experiments shown on this paper are all conducted on a single machine with the Intel(R) Xeon(R) CPU @ 2.20GHz model, the actual scheduling process for each semester may utilize various machines due to time constraints.
@@ 5.1 Performance of the Scheduling Models Across Multiple Semesters
Student Size Exam Size Exam Time Range # Available Time Slots Max-Clique Size
@@
To ensure a fair comparison, we use the full set of time slots for all models as shown in Table 1. For each semester, the number of blocks used remain consistent. For Layer-Cake-based models, the time slot configuration input matches that obtained from the corresponding GtS models. Unless otherwise specified, we use the standard scheduling parameters given in Appendix 12 throughout the paper. The Baseline model follows the historical approach described in Section 1. The current course schedule includes new meeting time patterns that are not accounted for in the lecture time-based mapping used in the baseline approach. To adapt the Baseline model for different semesters, we assigned courses with new meeting times greedily to one of the existing time slots. This approach preserved the direct mapping between meeting times and exam times.
@@ 5.2 Flexibility of the Framework
The results presented in the following subsections highlight the flexibility of our framework. These results are derived from routine experiments conducted each semester, allowing us to explore a variety of potential optimal schedules and present trade-offs to the University Registrar. Specifically, we focus on Spring 2024, the most recent semester at the time this paper is written.
@@ 5.2.2 Flexibility in Restricting Time Slot Availability
Table 4 presents the results of these exclusions. As more slots are restricted, the model’s options to optimize metrics diminishes. We observed a noticeable increase in the number of 2-in-24hrs and triples, suggesting that restricting multiple time slots can result in less optimal student schedules. In response, the team has been working with the Registrar to open additional time slots, which take effects starting from SP24.
@@
5.2.3 Flexibility in Weighing Different Metrics
Penalty weights are important parameters to conduct experiments on since they can significantly influence the balance between different metrics. We vary the weight of triples versus B2Bs in the objective of the Block_Sequencing_IP, adjusting the penalty for triples to be equivalent to a multiple of B2Bs, ranging from 1.5 to 15.5.
Figure 6 illustrates the resulting trade-off between B2Bs and triples on the GtS schedule. As the penalty weight for triples increases, the number of triples decreases, but this comes at the cost of an increased number of B2Bs. This demonstrates that prioritizing the reduction of triples leads to a higher occurrence of B2Bs, and vice versa.
Figure 6: Effect of Different B2Bs vs. Triples Weights on the GtS Schedule in Spring 2024
@@ 5.2.4 Flexibility in Front-Loading Large Exams
In this section, we explore the impact of different exam size and time slot cutoffs for front-loading large exams. We modify the Block Sequencing stage by experimenting with a range of size thresholds for large exams and corresponding early time slot cutoffs.
@@ 5.3 Regression Analysis on Key Input Parameters
In the previous section, we have observed that the number of blocks, metric weights, size cutoff, and slot cutoff can crucially impact the metrics. To better understand the trade-offs among these parameters, we generate schedules for every possible parameter combination (see Table 5) over three spring semesters. In total, we produce 2430 distinct schedules (810 per semester) and computed their metrics. We then fit a Random Forest Regression model on these results and perform a SHAP Value analysis (Lundberg and Lee 2017) to evaluate the influence of each parameter on the metrics. SHAP values provide a measure of feature importance, quantifying the contribution of each parameter to the model’s output. In this case, we choose the model output to be the number of students who would need rescheduled exams, as this is the most key concern. Recall that, at Cornell University, students with more than two exams in a 24-hour period are eligible for a rescheduled exam.
@@ 6 Conclusions
This paper presents a MIP-based optimization framework that leverages a multi-stage Group-then-Sequence approach combined with a Layer-Cake heuristic algorithm to automate the final exam scheduling process at Cornell University. Over five consecutive semesters, we have collaborated closely with the University Registrar to schedule final exams for the entire student body, consistently generating optimized schedules that outperform historical baselines. Our results demonstrate not only the effectiveness of our method but also its capacity to adapt to a wide range of constraints and objectives.

### v3

@@ 1 Introduction
1.1 Contributions
Our main contribution are as follows:
  1. 1.
  2. 2.
    We develop an integrated MIP-based framework that synergizes local search and heuristic techniques to efficiently generate exam schedules optimized according to multiple performance metrics. Moreover, our framework supports hybridization strategies to further enhance solution quality.
  3. 3.
1.2 Organization
The rest of the paper is organized as follows: First, we detail Cornell University’s final exam scheduling problem in Section 2. In Section 3, we provide a brief review of relevant work. Section 4 outlines our proposed solution methods, with detailed optimization models and algorithms provided in the Appendix. In Section 5, we present the outcomes of our scheduling models and algorithms using real student enrollment data from Cornell University. Finally, Section 6 concludes our work.
2 Problem Description
Cornell University faces the complex task of scheduling final exams for a significant student body of more than 15,000 and approximately 700 courses at the beginning of every semester. These exams are concentrated within a 7 to 9-day final exam period, with three time slots available per day (9am, 2pm, and 7pm). However, certain time slots—such as Saturday nights and Sunday mornings—may have restrictions, resulting in roughly 19 to 22 available time slots. The ultimate scheduling problem is to assign around 500 exams to these time slots. The physical locations for the exams are determined separately by the University Registrar, and as such, this study focuses solely on optimizing the timing of the exams. Room assignment is not a significant concern at Cornell University for two primary reasons. First, the university has a sufficient number of classrooms with adequate seating capacities to accommodate demand. The only potential constraint is a cap of 5,000 students taking exams per day; however, all generated schedules remain well below this limit. Second, Cornell University typically does not split exams across multiple classrooms, which further simplifies the assignment of physical locations.
The goal of an optimized final exam schedule is to enhance the experience for both students and faculty. From the faculty’s perspective, student conflicts—such as overlapping or consecutive exams–may require the scheduling of make-up exams according to university policies (https://registrar.cornell.edu/calendars-exams/final-exam-policies), which imposes additional burdens. Meanwhile, students desire a well-distributed exam schedule to allow for sufficient review and study time. As a result, the optimal exam schedule should minimize various unfavorable metrics, each with its own weight. These metrics include direct conflicts, back-to-back exams, two exams in 24 hours, three exams in a row, and three exams in 4 slots.
@@ 3.1 Educational Timetabling
Educational timetabling includes three main subfields: high-school timetabling, course timetabling, examination timetabling. Many real-world problems in these areas are approached through integer programming techniques. High-school timetabling, commonly represented in the XHSTT format (Post et al. 2016), has seen the application of MIP in studies like Kristiansen et al. (2015) and Fonseca et al. (2017). Course timetabling problems, such as those at the United States Air Force Academy (Gonzalez et al. 2018), the Technion in Israel (Strichman 2017), and MIT (Barnhart et al. 2022), also employ these approaches. In examination timetabling, there are case studies from the Universidad Politécnica de Madrid (García-Sánchez et al. 2019) and the Department of Education in Vestfold County (Avella et al. 2022)
@@ 4.1 Multi-Stage Group-then-Sequence (GtS) Model
The idea behind the first two stages is to initially group the exams into a smaller number of exam blocks and then assign each block to a time slot.
@@ 4.1.1 Block Assignment
The Block Assignment stage is formulated as a MIP problem (Min_Conflict_IP) that assigns each exam to one of a specified number of blocks, denoted by $k$. It is assumed that $k\leq|S|$, where $S:=\{0,1,\ldots,|S|-1\}$ is the set of available time slots. The objective of this grouping is to minimize the total number of pairwise co-enrollments (direct conflicts) among exams within the same block. Appendix 7.1 provides the Min_Conflict_IP formulation. It is important to note that, at this stage, each block represents a group of exams that will be scheduled simultaneously, but is not yet associated with a specific time slot.
@@
We also propose a variation of the Block Assignment problem, named Zero_Conflict_IP, that imposes a strict zero-conflict constraint. We assume a fixed linear ordering of exam blocks $10,1,\ldots,k-$ 1 and declare only the consecutive pairs $(b,b+1)$, for $b=0,\ldots,k-2$, to be neighboring (so block 00 and block $k-1$ are not neighbors). The model imposes a hard constraint forbidding any co‐enrolled exams from occupying the same block, and it minimizes the total number of co‐enrollments across all such consecutive block pairs. This objective draws an analogy to the Traveling Salesman Problem (TSP)—except that we optimize a Hamiltonian path rather than a cycle, reflecting the linear sequence of blocks. The Zero_Conflict_IP formulation is available in Appendix 7.2. Figure 2 shows a three‐block example together with its path‐objective value. Compared to the minimum-conflict formulation, Zero_Conflict_IP has a more constrained feasible space, which increases the likelihood of infeasible solutions.
The feasibility of having a zero-conflict block assignment depends mainly on the required number of blocks $k$. While one way to find a lower bound on $k$ is to enumerate all possible values and check the feasibility of the Zero_Conflict_IP problem, a smarter way is to solve for the Maximum Clique (Max_Clique_IP) problem. Appendix 8 contains the Max_Clique_IP formulation. The Max_Clique_IP involves finding a subset of vertices in $G$ that forms a complete sub-graph, where every two distinct vertices are adjacent. In our context, the size of the max clique corresponds to a lower bound on the number of blocks for which it is feasible to have no conflicts.
Figure 2: This is an example of a zero-conflict block assignment. The value on each edge represents the number of co-enrollments. Blocks 0 and 1, as well as Blocks 1 and 2, are considered neighboring blocks. Consequently, the path-objective value can be calculated as the sum of the highlighted values.
@@ 4.1.2 Block Sequencing
After assigning exams to the designated blocks, we then allocate these blocks to time slots by sequencing them in Block Sequencing stage (Block_Sequencing_IP). In the Block Assignment phase, we focus mainly on direct conflicts, but in Block Sequencing, we begin to account for other levels of undesirable metrics. Appendix 9 presents the Block_Sequencing_IP formulation.
@@ 4.1.3 Post-Processing Local Search Algorithm
Multiple optimal solutions with the same number of direct conflicts can be generated in the Block Assignment stage. However, the schedules derived from these assignments may vary in other metrics, as the Block Assignment model only focuses on direct conflicts. As a result, the block assignment obtained may not be optimal when considering all metrics comprehensively. To address this limitation, we develop a local search algorithm that takes an existing exam schedule and iteratively reschedules subsets of exams with a specified size $n$. The full pseudocode for the Post-Processing local search algorithm is available in Appendix 10.
In each iteration, the algorithm selects the top $n$ exams to reschedule by assessing the contribution of each exam to the number of bad metrics, normalized by the number of students taking the exam. For example, an exam where 30% of its students have an exam right before or after it would be prioritized for rescheduling compared to an exam with only 5% of students in a similar situation. This normalization effectively ranks exams by the density of “bad” events rather than their absolute counts, allowing the algorithm to focus on the intensity of issues experienced per student. Empirically, this approach provides a more effective signal for guiding the local search.
Once the subset is selected, the algorithm solves a Schedule_IP problem to move the exams to time slots that minimize the number of B2Bs and 2-in-24hrs, without increasing the number of direct conflicts. The process continues iteratively, with the algorithm using adaptive step sizes to adjust the starting index of the rolling window with size $n$. If no improvement is found after rescheduling, a larger step forward is taken to explore new solutions. Conversely, if an improvement is observed, a smaller step backward is adopted to fine-tune the schedule. This adaptive approach enables the algorithm to efficiently navigate the solution space, balancing exploration with refinement.
@@ 4.2 Layer-Cake Heuristic Algorithm
The Layer-Cake heuristic algorithm offers an alternative to the GtS model and is inspired by the Post-Processing algorithm. Unlike the GtS model, Layer-Cake addresses conflicts and other metrics simultaneously by solving the Schedule_IP problem, which optimizes all metrics together. To reduce the computational complexity of scheduling all exams at once, the problem is solved for a smaller subset of exams each time. A key difference between Layer-Cake and GtS is that Layer-Cake directly takes in a sequence of time slots, eliminating the need to specify the number of blocks, $k$, as an input. Appendix 11 provides the full algorithm pseudocode for Layer-Cake.
Figure 5: Visualization of the Layer-Cake Algorithm with Partial Overlaps in Consecutive Layers
Layer-Cake follows a layered approach, where each layer consists of a set of exams such that the total number of students per layer is similar. The layers are determined by sorting the exams in descending order of student enrollment. The initial layers focus primarily on larger exams, while subsequent layers include many smaller exams. The algorithm starts by scheduling the largest exams independently of the others. It then schedules the next subset of exams to build upon those already assigned, gradually building the schedule “layer by layer.” Crucially, each new layer can partially overlap with the preceding one, as depicted in Figure 5. This ordering ensures that the largest exams are scheduled first, as the Schedule_IP problem becomes more challenging when dealing with larger exams compared to smaller ones, even with a constant total number of students.
An essential feature of the Layer-Cake algorithm is the overlap between consecutive layers. In each iteration, a subset of exams that have already been scheduled are included in the IP, allowing them to be rescheduled alongside the newly introduced exams. This overlapping of layers, serving as a warm start for the IP problems, considerably boosts the algorithm’s performance.
@@ 5 Implementation Results and Impact
Once the team gains access to the necessary data, we conduct experiments using various parameters and models implemented in Gurobi and Python. This collaborative approach allows for comprehensive exploration and evaluation of different scheduling options. To accelerate the process, we use Google Colab to run multiple experiments in parallel. From the generated subset of “good” schedules, the University Registrar selects the best one to be published on the university website.
In the subsequent parts, we showcase the performance of our scheduling models on real student enrollment data, highlighting the framework’s ability to outperform the historical baseline and its flexibility in meeting the registrar’s needs. All experiments in this paper are conducted on a single machine equipped with an 8-core Intel Xeon CPU running at 2.20 GHz. All optimization models are solved using Gurobi version 11.0.3. Computation times are not reported in this paper, as they are not a primary concern in our application. All methods complete within several hours, which is acceptable given that the final exam schedule is typically prepared over the span of approximately one month each semester.
@@ 5.1 Performance of the Scheduling Models Across Multiple Semesters
Student Size # of Unique Exams Exam Time Range # of Available Time Slots Max-Clique Size
@@
To ensure a fair comparison, we use the full set of time slots for all models as shown in Table 1. For each semester, the number of blocks used remains consistent across different scheduling model variants. For Layer-Cake-based models, the time slot configuration input matches that obtained from the corresponding GtS models. Unless otherwise specified, we use the standard scheduling parameters given in Appendix 12 throughout the paper. The Baseline model follows the historical approach described in Section 1. The current course schedule includes new meeting time patterns that are not accounted for in the lecture time-based mapping used in the baseline approach. To adapt the Baseline model for different semesters, we assigned courses with new meeting times greedily to one of the existing time slots. This approach preserved the direct mapping between meeting times and exam times.
@@ 5.2 Flexibility of the Framework
The results presented in the following subsections highlight the flexibility of our framework. These results are derived from routine experiments conducted each semester, allowing us to explore a variety of high-quality schedules and present trade-offs to the University Registrar. Specifically, we focus on Spring 2024, the most recent semester at the time this paper is written.
@@ 5.2.2 Flexibility in Restricting Time Slot Availability
Table 4 presents the results of these exclusions. As more slots are restricted, the model’s options to optimize metrics diminishes. We observed a noticeable increase in the number of 2-in-24hrs and triples, suggesting that restricting multiple time slots can result in less favorable student schedules. In response, the team has been working with the Registrar to open additional time slots, which take effects starting from SP24.
@@ 5.2.3 Flexibility in Front-Loading Large Exams
In this section, we explore the impact of different exam size and time slot cutoffs for front-loading large exams. We modify the Block Sequencing stage of the GtSP model by experimenting with a range of size thresholds for large exams and the corresponding early time slot cutoffs.
@@ 5.3 Regression Analysis on Key Input Parameters
In the previous section, we have observed that the number of blocks, metric weights, size cutoff, and slot cutoff can crucially impact the metrics. To better understand the trade-offs among these parameters, we use the GtSP model to generate schedules for all combinations of parameter settings listed in Table 5, evaluated over the three spring semesters. In total, we produce 2430 distinct schedules (810 per semester) and computed their metrics. We then fit a Random Forest Regression model on these results and perform a SHAP Value analysis (Lundberg and Lee 2017) to evaluate the influence of each parameter on the metrics. SHAP values provide a measure of feature importance, quantifying the contribution of each parameter to the model’s output. In this case, we choose the model output to be the number of students who would need rescheduled exams, as this is the most key concern. Recall that, at Cornell University, students with more than two exams in a 24-hour period are eligible for a rescheduled exam.
@@ 6 Conclusions
This paper presents a MIP-based optimization framework that leverages a multi-stage Group-then-Sequence approach combined with a Layer-Cake heuristic algorithm to automate the final exam scheduling process at Cornell University. Over five consecutive semesters, we have collaborated closely with the University Registrar to schedule final exams for the entire student body, consistently generating optimized schedules that outperform historical baselines. Our results demonstrate not only the effectiveness of our method but also its capacity to adapt to a wide range of constraints and objectives. Further evidence of the framework’s generalizability is provided in Appendix 13, which demonstrates its competitive performance against established exam timetabling algorithms on the University of Nottingham dataset.